# A Modular Approach for Multimodal Summarization of TV Shows

### v3

@@ Abstract
In this paper we address the task of summarizing television shows, which touches key areas in AI research: complex reasoning, multiple modalities, and long narratives. We present a modular approach where separate components perform specialized sub-tasks which we argue affords greater flexibility compared to end-to-end methods. Our modules involve detecting scene boundaries, reordering scenes so as to minimize the number of cuts between different events, converting visual information to text, summarizing the dialogue in each scene, and fusing the scene summaries into a final summary for the entire episode. We also present a new metric, PREFS (Precision and Recall Evaluation of Summary Facts), to measure both precision and recall of generated summaries, which we decompose into atomic facts. Tested on the recently released SummScreen3D dataset Papalampidi and Lapata (2023), our method produces higher quality summaries than comparison models, as measured with ROUGE and our new fact-based metric.
@@ 1 Introduction
Most prior work on creative summarization does not consider the above challenges all at once, focusing either on the text modality and full-length narratives with complex semantics Gorinski and Lapata (2015); Chen et al. (2022); Agarwal et al. (2022) or on short video clips which last only a couple of minutes Tapaswi et al. (2016); Lei et al. (2018); Liu et al. (2020). A notable exception is Papalampidi and Lapata (2023), who incorporate multimodal information into a pre-trained textual summarizer by adding (and tuning) adapter layers Rebuffi et al. (2017); Houlsby et al. (2019). On the evaluation front, there is no single agreed-upon metric for measuring summary quality automatically, although there is mounting evidence that ROUGE Lin (2004) does not discriminate between different types of errors, in particular those relating to factuality Min et al. (2023); Clark et al. (2023).
@@
As scene breaks are not always given explicitly, we devise an algorithm to identify them from the order of the speaker names (row 1, Figure 1). Additionally, we select the optimal order in which to re-arrange scenes (row 2, Figure 1), as these often appear in non-linear order (e.g., , cuts between different subplots or flashbacks). Next, we produce summaries in a two-layer process. A vision-processing module Lei et al. (2020); Lin et al. (2022) converts the video to text using visual captioning (row 3, Figure 1), which leverages the strong specialized ability of vision-to-text models and allows us to treat the problem as one of text-to-text summarization. We also summarize each scene independently with a module specialized for dialogue summarization (row 4, Figure 1). Finally, we use a high-level summarization module specialized for narrative summarization to fuse the sequence of scene summaries into a final summary (row 5, Figure 1).
We also propose a new metric for assessing the factuality of generated summaries by adapting FActScore Min et al. (2023), a recently introduced metric for detecting hallucination in text generation. We break the generated summary into atomic facts, and check what fraction of them are supported by the reference. This we term fact-precision. We also do the same in reverse, breaking the reference into facts and measuring what fraction are supported by the generated summary, which we term fact-recall. Our metric, PREFS (Precision and Recall Evaluation of Summary Facts), is the harmonic average of these two scores. Our contributions are:
@@ 2 Related Work
3 Decomposition of Multimodal Summarization into Modules
@@ 3.1 The Multimodal Summarization Task
Before discussing the details of the various modules, we provide information on our specific task and the dataset we are working with. We develop and evaluate our approach on SummScreen3D Papalampidi and Lapata (2023) [^1] https://github.com/ppapalampidi/video_abstractive_summarization, which to our knowledge is the only existing dataset for long-form video summarization. It consists of 5,421 videos of TV episodes (mostly soap operas) varying in length from 30–60min, with accompanying transcripts (on average 6K tokens long) and summaries that were written by fans and scraped from public websites (average length is 200 tokens). These are partitioned into 296 each for validation and testing, and the remaining 4,829 for training. Videos can have multiple summaries from different fansites (the average number of summaries per episode is 1.53), giving a total of 8,880 training pairs.
In SummScreen3D, each data point contains a video, a transcript that includes character names and, sometimes, marked scene breaks, and a closed captions file, which is used to display subtitles and has timestamps but not speaker names. We align the lines in the transcript with those from the closed captions. These do not match perfectly, because of slight errors in the automatic transcription. For each line $t$ in the transcript and utterance $c$ in the caption, we estimate a similarity score as $\tfrac{|f(t,c)|}{\min{(|t|,|c|)}}$, where $f$ computes longest common subsequence. Then we use dynamic time warping to align both sequences Myers and Rabiner (1981); Papalampidi et al. (2021). As a result, we obtain alignments of transcript utterances to video segments, which allows us to use the scene-breaks from the former to segment the latter.
@@ 3.2 Scene Detection
We define a cost for a given partition, by invoking the minimum description length (MDL) principle Grünwald (2007), in which the optimal representation of a piece of data is that which contains the fewest of bits. MDL has been shown to have many nice theoretical properties (Grünwald, 2007) and has been used in recent machine learning applications from interpreting hidden representations Voita and Titov (2020) to measuring noise-free information content in images Mahon and Lukasiewicz (2024). We use it here to formalize the desirable properties of a good segmentation.
Thus, the cost of a partition is the number of bits needed to specify it. We assume the total set of $N$ character names for the entire transcript is given. Then, for each scene, we make a scene-specific codebook for the $n$ characters that appear there, which assigns each character a code as an index from $10,\dots,n-$ 1, which all have length $\leq\lceil\log{n}\rceil$. Using exactly $\lceil\log{n}\rceil$ guarantees a prefix-free code Grunwald (1998) [^2] Because we do not need to convert to concrete codes, we omit the ceiling operator in our optimization objective.. The number of possible codebooks with $n$ out of $N$ characters is ($N\choose n$. Imposing e.g., lexicographic order, on these, the number of bits to specify one is
@@
where $N$ is the total number of speakers as before, and $n$ and $l$ are, respectively, the number of distinct speakers and the number of transcript lines in the scene. See Appendix D for details and examples. Intuitively, the two terms in (1) impose opposite pressures on the solution. Term 1, the bitcost of specifying a new code book, imposes a cost on each scene break, and means the algorithm only inserts a scene break if it leads to a sufficient reduction in term two to outweigh this cost. Term 2, the bitcost of using each code book to specify the characters in the corresponding scene, penalizes solutions in which scenes contain lots of different characters, and encourages instead segmentations in which the number of speaking characters in each scene is minimized.
@@
Our algorithm runs in $O(N^{3})$, with only $O(N^{2})$ calls to compute the scene cost as per Equation (1). The scene breaks can be transferred to the video because of the dynamic time-warping alignment between the transcript and the timestamps in the closed captions, as described in Section 3.1. Note, we do not have to specify the number of scenes, that is determined automatically by our algorithm.
@@ 3.5 Summary Generation
Because of our two-layer summarization approach, we can use a relatively small backbone model, and are still able to encode very long input. In our experiments, we use variants of BART-large Lewis et al. (2020) for both the dialogue summarization and high-level summarization modules, but any other similar model could be used instead. For the dialogue summarization module (Figure 1, row 4), we use the public Huggingface checkpoint for BART which has been fine-tuned on the SamSum dataset Gliwa et al. (2019), to output multi-sentence summaries of dialogue. For the high-level summarization module (Figure 1, row 5), we use BART, fine-tuned first for document summarization on the CNN/Daily Mail dataset Hermann et al. (2015), and then on SummScreen3D. The input for the latter fine-tuning is our re-ordered scene summaries and visual text descriptions, i.e., the output of the dialogue summarization, vision-processing and reordering modules. The output is the gold summaries from the SummScreen3D Papalampidi and Lapata (2023) training set. Training and inference took place on a single NVIDIA A100-SXM-80GB GPU, taking seven and one hour(s), respectively.
4 Fact-Precision, Fact-Recall and PREFS
@@
We adapt this for summary evaluation by replacing the external knowledge source with the gold summaries. Note that this incorporates relevance as well as accuracy. There are many true facts about the TV show being summarized, and only some of them appear in the gold summaries. We assume that facts that do not appear in the gold summary are irrelevant and so should not be marked as correct if they appear in our model summaries. This is our fact-precision metric. For fact-recall, we do the same in reverse: we convert the gold summary into atomic facts, and then check whether each of these is supported by the generated summary. Again, the score is the fraction of such facts that are supported. A summary will get a high fact-precision score if every atomic piece of information is both true and relevant enough to appear in the gold summary. It will get a high fact-recall score if it contains every atomic piece of information that was contained in the gold summary. The final score for our metric, which we term PREFS (Precision and Recall Evaluation of Summary Facts), is the harmonic mean of fact-precision $FP$ and fact-recall $FR$:
$\operatorname{PREFS}=\frac{2}{\frac{1}{FP}+\frac{1}{FR}}\,.$
@@
We examined whether PREFS correlates with human judgments of factuality. Two annotators, both English native speakers, were asked to watch four randomly selected shows from the SummScreen3D development set. They were then shown facts generated by GPT4 corresponding to summaries produced by several automatic systems (those described in Section 5.1). For each fact they were asked to decide whether it was supported by the episode (1/True or 0/False). Human ratings significantly correlate with GPT4’s estimate on whether a fact is supported (Pearson’s $r=0.5$, $N=520$, $p<0.01$), with an inter-annotator agreement of $r=0.57$ ($p<0.01$) as upper bound.
@@ 5 Experimental Evaluation
r1 r2 rlsum fact-prec fact-rec PREFS
@@
Table 1: Automatic evaluation results on SummScreen3D. The first block presents text-only models and the second one multimodal ones. Best results are in bold, second-best in italics. For ROUGE, we report the mean of five independent random seeds, with standard deviation in parentheses.
@@ 5.1 Comparison Models
Our full model was composed of the modules described in Section 3. The scene detection module was employed on SummScreen3D transcripts without explicitly marked breaks (approximately 20% of the time). The high-level summarization module was trained for a max of 10 epochs, with early stopping using ROUGE on the validation set, with a patience of 2. The optimizer was AdamW with learning rate 1e-6, and a linear scheduler with 0 warmup steps. The other models in our framework are not fine-tuned.
We compare our approach to the end-to-end model of Papalampidi and Lapata (2023), which, to our knowledge, represents the state of the art on our task, and various text-only models which operate on the transcript:
@@
Bertsch et al. (2023) is a retrieval-based method that is particularly suited to processing long inputs. It wraps around any transformer, with context size $k$, and can extend this size arbitrarily by storing an index of all input tokens and replacing the transformer query mechanism with the $k$-nearest neighbours from this index.
@@
r1 r2 rlsum fact-prec fact-rec PREFS
@@
Table 2: Effect of removing various modules from our method.
@@ 5.2 Results
Table 1 summarizes our results, as measured by ROUGE-1 (r1), ROUGE-2 (r2), ROUGE-Lsum (rlsum) (computed using the python-rouge package) and our new PREFS metric. We present results with BERTScore Zhang et al. (2019) in Appendix E, for the sake of brevity. Results for Papalampidi and Lapata (2023), which we denote as ‘adapter-e2e’, were reproduced with their code. [^4] https://github.com/ppapalampidi/video_abstractive_summarization For PREFS, as described in Section 4, we use a separate query to GPT for each fact to check whether it is supported. There are roughly 70 facts per generated summary, which can lead to significant financial cost if used excessively. Therefore, while we report five random seeds for ROUGE, we select a single seed (randomly) for fact-precision and fact-recall. Example output is shown in Appendix A.
@@
PREFS better reflects summary quality than ROUGE.
@@
Interestingly, upper-bound ROUGE scores are lower than those for several models. We regard this as a shortcoming of ROUGE as a metric. Qualitatively, reading the different gold summaries shows that they are more similar to and accurate with respect to each other than any of the predicted summaries, including ours. Although they differ in phrasing and length, they all describe the same key events. Our proposed metric, PREFS, better reflects this similarity and gives a much higher score to ‘upper-bound’ than to any of the models. The large gap to get to the level of ‘gold upper-bound’ reflects the difficulty of the task. The comparison between ROUGE and PREFS suggests that the former is useful for detecting which summaries are of very low quality, e.g., if ROUGE-2 is $<5$, the summary can confidently be regarded as poor. However, for distinguishing between multiple relatively decent summaries, PREFS is more useful.
Ablations show all modules are important, but the transcript is the most important of all.
@@
Table 2 shows the effect of removing four of the five modules. In ‘w/o scene-detection’, we remove the scene-detection module and just split the input into equally-sized chunks equal to the context size (1,024). In ‘w/o transcript’ we do not include summaries of the dialogue from the transcript, and the only input to the high-level summarization module is the output of the visual processing module. In ‘w/o video’ we do the reverse: use only the dialogue summaries without the output of the vision-processing module. In ‘w/o reordering’, we remove the scene-reordering module and present the scene summaries to the high-level summarization module in the order in which they appear. All these experiments are performed using the Kosmos-2 captions (except ‘w/o video’ which has no captions). As in Table 1, ROUGE is scored over five random seeds, PREFS for a single randomly selected seed.
@@
Aside from ROUGE-1 and ROUGE-Lsum, which are roughly the same for ‘w/o reordering’ and ‘w/o scene-detection’, all ablation settings lead to a drop in all metrics. We provide detailed analysis of the accuracy of the scene detection module in Appendix 5.3. Unsurprisingly, the largest drop is in ‘w/o transcript’ as most information is in the dialogue — a TV show without sound or subtitles is difficult to follow. Interestingly, however, this setting still gets some n-grams (ROUGE) and facts (PREFS) correct, showing that our model is extracting useful information from the video. This is also clear from the ‘w/o video’ setting, which shows a drop in all metrics. Qualitatively, we observe that the vision-processing module is most useful for identifying locations, e.g., ‘at the hospital’ which is generally not mentioned in the dialogue. For ‘w/o scene-detection’ and ‘w/o reordering’, the difference is moderate for ROUGE. For PREFS, it is more substantial and highly significant when taken over all facts in the dataset.
5.3 Accuracy of Scene-detection Algorithm
@@
Here, in Table 3, we report an empirical test of the accuracy of our scene-detection algorithm. These figures are produced using an episode which has explicitly marked scene breaks. We assign each transcript line a label based on the scene it is in with respect to these explicit scene breaks, e.g., all lines in the first scene get the label ‘0’. Then we do the same for the scene breaks produced by our algorithm, and compare the two sets of labels using unsupervised label comparison metrics, as commonly used in clustering problems: accuracy (ACC), normalized mutual information (NMI), and adjusted Rand index (ARI), defined as, e.g., in Mahon and Lukasiewicz (2023). We compare to two baselines. The first, denoted ‘uniform’, divides each episode into $n$ equally-sized scenes, where $n$ is the average number of scenes per episode in the dataset. The second, denoted ‘uniform oracle’, does the same except sets $n$ to the true number of scenes in that episode.
@@
Our method produces more accurate scene splits than both baselines, despite taking no information from the ground-truth labels. Moreover, many of the occasions on which it differs from the ground-truth scene splits appear at least as reasonable as the ground-truth splits when we observe only the character names in the transcript. This suggests that errors in the predicted splits are due not to the algorithm itself but to the fact that, at present, it only uses character names and ignores the speech itself. A future extension is to use named entities or all nouns from the speech as well as character names.
@@
Table 3: The accuracy of the scene splits produced by our method, benchmarked against two methods that split the transcript uniformly: ‘uniform’ splits into the average number of scenes in the dataset; ‘uniform oracle’ into the ground-truth number of scenes for each episode.

### v4

@@ Abstract
In this paper we address the task of summarizing television shows, which touches key areas in AI research: complex reasoning, multiple modalities, and long narratives. We present a modular approach where separate components perform specialized sub-tasks which we argue affords greater flexibility compared to end-to-end methods. Our modules involve detecting scene boundaries, reordering scenes so as to minimize the number of cuts between different events, converting visual information to text, summarizing the dialogue in each scene, and fusing the scene summaries into a final summary for the entire episode. We also present a new metric, PRisma (Precision and Recall Evaluation of Summary Facts), to measure both precision and recall of generated summaries, which we decompose into atomic facts. Tested on the recently released SummScreen3D dataset Papalampidi and Lapata (2023), our method produces higher quality summaries than comparison models, as measured with ROUGE and our new fact-based metric.
@@ 1 Introduction
Most prior work on creative summarization does not consider the above challenges all at once, focusing either on the text modality and full-length narratives with complex semantics Gorinski and Lapata (2015); Chen et al. (2022); Agarwal et al. (2022) or on short video clips which last only a couple of minutes Tapaswi et al. (2016); Lei et al. (2018); Liu et al. (2020). A notable exception are Papalampidi and Lapata (2023), who incorporate multimodal information into a pre-trained textual summarizer by adding (and tuning) adapter layers Rebuffi et al. (2017); Houlsby et al. (2019). On the evaluation front, there is no single agreed-upon metric for measuring summary quality automatically, although there is mounting evidence that ROUGE Lin (2004) does not discriminate between different types of errors, in particular those relating to factuality Min et al. (2023); Clark et al. (2023).
@@
As scene breaks are not always given explicitly, we devise an algorithm to identify them from the order of the speaker names (row 1, Figure 1). Additionally, we select the optimal order in which to re-arrange scenes (row 2, Figure 1), as these often appear non-linearly (e.g., cuts between different subplots or flashbacks). Next, we produce summaries in a two-layer process. A vision-processing module Lei et al. (2020); Lin et al. (2022) converts the video to text using visual captioning (row 3, Figure 1), which leverages the strong specialized ability of vision-to-text models and allows us to treat the problem as one of text-to-text summarization. We also summarize each transcript scene independently with a module specialized for dialogue summarization (row 4, Figure 1). Finally, a high-level summarization module specialized for narrative summarization fuses the sequence of scene summaries into a final summary (row 5, Figure 1).
We also propose a new metric for assessing the factuality of generated summaries by adapting FActScore Min et al. (2023), a recently introduced metric for detecting hallucination in text generation. We break the generated summary into atomic facts, and check what fraction of them are supported by the reference. This we term fact-precision. We also do the same in reverse, breaking the reference into facts and measuring what fraction are supported by the generated summary, which we term fact-recall. Our metric, PRisma (Precision and Recall Evaluation of Summary Facts), is the harmonic average of these two scores. Our contributions are:
@@ 2 Related Work
3 Decomposition into Modules
@@ 3.1 The Multimodal Summarization Task
Before discussing the details of the various modules, we provide information on our specific task and the dataset we are working with. We develop and evaluate our approach on SummScreen3D [^1] https://github.com/ppapalampidi/video_abstractive_summarization Papalampidi and Lapata (2023), which to our knowledge is the only existing dataset for long-form video summarization. It consists of 5,421 videos of TV episodes (mostly soap operas) varying in length from 30–60min, with accompanying transcripts (on average 6K tokens long) and summaries that were written by fans and scraped from public websites (average length is 200 tokens). These are partitioned into 296 each for validation and testing, and the remaining 4,829 for training. Videos can have multiple summaries from different fansites (the average number of summaries per episode is 1.53), giving a total of 8,880 training pairs.
In SummScreen3D, each data point contains a video, a transcript that includes character names and, sometimes, marked scene breaks, and a closed captions file, which is used to display subtitles and has timestamps but not speaker names. We align the lines in the transcript with those from the closed captions. These do not match perfectly, because of slight errors in the automatic transcription. For each line $t$ in the transcript and utterance $c$ in the caption, we estimate a similarity score as $\tfrac{|f(t,c)|}{\min{(|t|,|c|)}}$, where $f$ computes the longest common subsequence. Then we use dynamic time warping to align both sequences Myers and Rabiner (1981); Papalampidi et al. (2021). As a result, we obtain alignments of transcript utterances to video segments, which allows us to use the scene-breaks from the former to segment the latter.
@@ 3.2 Scene Detection
We define a cost for a given partition, by invoking the minimum description length (MDL) principle Grünwald (2007), in which the optimal representation of a piece of data is that which contains the fewest of bits. MDL has been shown to have many nice theoretical properties (Grünwald, 2007) and has been used in recent machine learning applications from interpreting hidden representations Voita and Titov (2020) to measuring noise-free information content in images (Mahon and Lukasiewicz, 2024). We use it here to formalize the desirable properties of a good segmentation.
Thus, the cost of a partition is the number of bits needed to specify it. We assume the total set of $N$ character names for the entire transcript is given. Then, for each scene, we make a scene-specific codebook for the $n$ characters that appear there, which assigns each character a code as an index from $10,\dots,n-$ 1, which all have length $\leq\lceil\log{n}\rceil$. Using exactly $\lceil\log{n}\rceil$ guarantees a prefix-free code Grunwald (1998) [^2] Because we do not need to convert to concrete codes, we omit the ceiling operator in our optimization objective.. The number of possible codebooks with $n$ out of $N$ characters is ($N\choose n$. Imposing some order, e.g., lexicographic, on these, the number of bits to specify one is
@@
where $N$ is the total number of speakers as before, and $n$ and $l$ are, respectively, the number of distinct speakers and the number of transcript lines in the scene. See Appendix D for details and examples. Intuitively, the two terms in (1) impose opposite pressures on the solution. Term 1, the bitcost of specifying a new code book, imposes a cost on each scene break, and means the algorithm only inserts a scene break if it leads to a sufficient reduction in term two to outweigh this cost. Term 2, the bitcost of using each code book to specify the characters in the corresponding scene, penalizes solutions where scenes contain lots of different characters, and encourages instead segmentations where the number of speaking characters in each scene is minimized.
@@
Our algorithm runs in $O(N^{3})$, with only $O(N^{2})$ calls to compute the scene cost as per Equation (1). The scene breaks can be transferred to the video because of the dynamic time-warping alignment between the transcript and the timestamps in the closed captions, as described in Section 3.1. Note, that our algorithm determines automatically the number of scenes.
@@ 3.5 Summary Generation
Because of our two-layer summarization approach, we can use a relatively small backbone model, and are still able to encode very long input. In our experiments, we use variants of BART-large Lewis et al. (2020) for both the dialogue summarization and high-level summarization modules, but any other similar model could be used instead. For the dialogue summarization module (Figure 1, row 4), we use the public Huggingface checkpoint for BART which has been fine-tuned on the SamSum dataset Gliwa et al. (2019), to output multi-sentence summaries of dialogue. For the high-level summarization module (Figure 1, row 5), we use BART, fine-tuned first for document summarization on the CNN/Daily Mail dataset Hermann et al. (2015), and then on SummScreen3D. The input for the latter fine-tuning is our re-ordered scene summaries and visual text descriptions, i.e., the output of the dialogue summarization, vision-processing and reordering modules. The output is the gold summaries from the SummScreen3D Papalampidi and Lapata (2023) training set.
4 Fact Precision, Fact Recall and PRisma
@@
We adapt this for summary evaluation by replacing the external knowledge source with the gold summaries. Note that this incorporates relevance as well as accuracy. There are many true facts about the TV show being summarized, and only some of them appear in the gold summaries. We assume that facts that do not appear in the gold summary are irrelevant and so should not be marked as correct if they appear in our model summaries. This is our fact-precision metric. For fact-recall, we do the same in reverse: we convert the gold summary into atomic facts, and then check whether each of these is supported by the generated summary. Again, the score is the fraction of such facts that are supported. A summary will get a high fact-precision score if every atomic piece of information is both true and relevant enough to appear in the gold summary. It will get a high fact-recall score if it contains every atomic piece of information that was contained in the gold summary. The final score for our metric, which we term PRisma (Precision and Recall Evaluation of Summary Facts), is the harmonic mean of fact-precision $FP$ and fact-recall $FR$:
$\operatorname{\textsc{PRisma}}=\frac{2}{\frac{1}{FP}+\frac{1}{FR}}\,.$
@@
We examined whether PRisma correlates with human judgments of factuality. Two annotators, both English native speakers, were asked to watch four randomly selected shows from the SummScreen3D development set. They were then shown facts generated by GPT4 corresponding to summaries produced by several automatic systems (those described in Section 5.1). For each fact they were asked to decide whether it was supported by the episode (1/True or 0/False). Human ratings significantly correlate with GPT4’s estimate on whether a fact is supported (Pearson’s $r=0.5$, $N=520$, $p<0.01$), with an inter-annotator agreement of $r=0.57$ ($p<0.01$) as upper bound.
@@ 5 Experimental Evaluation
r1 r2 rlsum fact-prec fact-rec PRisma
@@
Table 1: Automatic evaluation results on SummScreen3D. The first block presents text-only models and the second one multimodal ones. Best results are in bold, second-best in italics. For ROUGE, we report the mean of five independent random seeds, with standard deviation in parentheses. We present example output in Appendix A.
Our full model was composed of the modules described in Section 3. For those SummScreen3D transcripts without explicitly marked breaks (approximately 20%), we split into scenes using Algorithm 1. The high-level summarization module was trained for a max of 10 epochs, with early stopping using ROUGE on the validation set, with a patience of 2. The optimizer was AdamW with learning rate 1e-6, and a linear scheduler with 0 warmup steps. Training and inference took place on a single NVIDIA A100-SXM-80GB GPU, taking seven and one hour(s), respectively. The other models in our framework are not fine-tuned.
@@ 5.1 Comparison Models
We compare our approach to the end-to-end model of Papalampidi and Lapata (2023), which represents the state of the art on our task, and various text-only models operating on the transcript:
@@
Bertsch et al. (2023) is a retrieval-based method that is particularly suited to processing long inputs. It wraps around encoder-decoder transformers, with context size $k$, and can extend this size arbitrarily by storing an index of all input tokens and replacing the cross-attention query mechanism with the $k$-nearest neighbours from this index.
@@
r1 r2 rlsum fact-prec fact-rec PRisma
@@
Table 2: Effect of removing various modules from our method (modular-kosmos).
@@ 5.2 Results
Table 1 summarizes our results, as measured by ROUGE-1 (r1), ROUGE-2 (r2), ROUGE-Lsum (rlsum) (computed using the python-rouge package) and our new PRisma metric. We present results with BERTScore Zhang et al. (2019) in Appendix E, for the sake of brevity. Results for Papalampidi and Lapata (2023), which we denote as ‘adapter-e2e’, were reproduced with their code. [^4] https://github.com/ppapalampidi/video_abstractive_summarization For PRisma, we use a separate query to GPT for each fact to check whether it is supported. There are roughly 70 facts per generated summary, which can lead to significant financial cost if used excessively. Therefore, while we report five random seeds for ROUGE, we select a single seed (randomly) for fact-precision and fact-recall.
@@
PRisma better reflects summary quality than ROUGE.
@@
Interestingly, upper-bound ROUGE scores are lower than those for several models. We regard this as a shortcoming of ROUGE as a metric. Qualitatively, reading the different gold summaries shows that they are more similar to and accurate with respect to each other than any of the predicted summaries, including ours. Although they differ in phrasing and length, they all describe the same key events. PRisma better reflects this similarity and gives a much higher score to ‘upper-bound’ than to any of the models. The large gap to get to the level of ‘gold upper-bound’ reflects the difficulty of the task. The comparison between ROUGE and PRisma suggests that the former is useful for detecting which summaries are of very low quality, (e.g., if ROUGE-2 is $<5$). However, for distinguishing between multiple relatively decent summaries, PRisma is more useful.
All modules are important, but the transcript is the most important of all.
@@
Table 2 shows the effect of removing four of the five modules. In ‘w/o scene-detection’, we remove the scene-detection module and just split the input into equally-sized chunks equal to the context size (1,024). In ‘w/o transcript’ we do not include summaries of the dialogue from the transcript, and the only input to the high-level summarization module is the output of the visual processing module. In ‘w/o video’ we do the reverse: use only the dialogue summaries without the output of the vision-processing module. In ‘w/o reordering’, we remove the scene-reordering module and present the scene summaries to the high-level summarization module in the order in which they appear. All these experiments are performed using the Kosmos-2 captions (except ‘w/o video’ which has no captions).
@@
Aside from ROUGE-1 and ROUGE-Lsum, which are roughly the same for ‘w/o reordering’ and ‘w/o scene-detection’, all ablation settings lead to a drop in all metrics. Unsurprisingly, the largest drop is in ‘w/o transcript’ as most information is in the dialogue — a TV show without sound or subtitles is difficult to follow. Interestingly, however, this setting still gets some n-grams (ROUGE) and facts (PRisma) correct, showing that our model is extracting useful information from the video. This is also clear from the ‘w/o video’ setting, which shows a drop in all metrics. Qualitatively, we observe that the vision-processing module is most useful for identifying locations, e.g., ‘at the hospital’ which is generally not mentioned in the dialogue. For ‘w/o scene-detection’ and ‘w/o reordering’, the difference is moderate for ROUGE. For PRisma, it is more substantial and highly significant when taken over all facts in the dataset.
Our scene-detection algorithm is more accurate than uniform baselines.
@@
Here, in Table 3, we report an empirical test of the accuracy of our scene-detection algorithm (Section 3.2, Algorithm 1). These figures are produced using an episode which has explicitly marked scene breaks. We assign each transcript line a label based on the scene it is in with respect to these explicit scene breaks, e.g., all lines in the first scene get the label ‘0’. Then we do the same for the scene breaks produced by our algorithm, and compare the two sets of labels using unsupervised label comparison metrics, as commonly used in clustering problems: accuracy (ACC), normalized mutual information (NMI), and adjusted Rand index (ARI), defined as, e.g., in Mahon and Lukasiewicz (2023). We compare to two baselines. The first, denoted ‘uniform’, divides each episode into $n$ equally-sized scenes, where $n$ is the average number of scenes per episode. The second, denoted ‘uniform oracle’, does the same except sets $n$ to the true number of scenes in that episode.
@@
Our method produces more accurate scene splits than both baselines, despite taking no information from the ground-truth labels. Moreover, many of the occasions on which it differs from the ground-truth scene splits appear at least as reasonable as the ground-truth splits when we observe only the character names in the transcript. This suggests that errors in the splits are not due to the algorithm itself but to the fact that, at present, it only uses character names, ignoring the speech itself. A future extension is to use named entities or all nouns from the speech as well as character names.
@@
Table 3: Accuracy of scene splits produced by our method, benchmarked against: ‘uniform’ splits into average number of scenes in the dataset; ‘uniform oracle’ into ground-truth number of scenes for each episode.
Humans prefer summaries generated by our modular approach
We selected 8 episodes from SummScreen3D, and asked human participants (recruited through the Prolific platform) to watch the episode, then rank summaries in order of perceived quality. The summaries we included in this study were ours, the gold summary, unlimiformer Bertsch et al. (2023), ‘adapter e2e‘ Papalampidi and Lapata (2023), and finetuned Mistral. We instructed participants to priortize informativeness and factual correctness, with a secondary emphasis on fluent well-written text. We included an attention-check question that participants would only be able to answer if they had watched the full episode, and asked for a short justification for the chosen ranking. The full instructions can be found in Appendix H. We recruited five annotators for each of the eight episodes.
Table 4 presents the ranking given to the systems by our participants. Specifically, we calculate Elo (1967) scores for each of the summary methods we perform evaluation on. We adjust scores once after all comparisons are made so as to be independent of the order of episodes. We are interested only in relative scores and correlations with other metrics, which are independent of the adjustment factor $K$ and the starting score. So, we set $K=1$ and the starting score to $\tfrac{1}{2}$, and the score for each method reduces to the fraction of pairwise comparisons they won. We can also then observe that our PRisma metric is a much better predictor of human preference (Pearson’s $r=0.792$) than ROUGE-2 ($r=0.594$) and ROUGE-Lsum ($r=0.512$).
Models elo
upper-bound 0.750
modular-kosmos 0.681
mistral-7b 0.388
adapter-e2e 0.353
unlimiformer 0.319
Table 4: Ranking given to summarization systems by human subjects; Elo scores from pairwise comparisons.